I would like some help with this problem:
In a new community, the sales representative tells you that there are 38 homes being built, with two floor plans. Floor Plan#1 sells for $175,000 and floor Plan #2 sells for $200,000. She mentions that all the available houses combined are worth $7,200,000. Write an equation that illustrates the solution, also use elimination to determine how many houses with each floor plan are available.
Answer by stanbon(75887) 
 In a new community, the sales representative tells you that there are 38 homes being built, with two floor plans. Floor Plan#1 sells for $175,000 and floor Plan #2 sells for $200,000. She mentions that all the available houses combined are worth $7,200,000. Write an equation that illustrates the solution, also use elimination to determine how many houses with each floor plan are available
----
Equations:
Quantity Eq.: x + y = 38
Value Eq:::175000x+200000y = 7200000
---------------------------------------------
Modify the 2nd equation:
175x + 200y = 7200
-------------
Multiply thru the 1st equation by 175 to get:
175x + 175y = 175*38
-------------------------
Subtract that from the equation above to get:
25y = 550
y = 22 (# of $200,000 homes)
----
Since x + y = 38, x = 16 (# of $175,000 homes)
==================================================
Cheers,
Stan H.
| |
|